OpenStudy (anonymous):

Find the y intercept of y=0.50(1.06)^x

mathslover (mathslover):

remember that whenever you are told to find the y intercept then you have to put x = 0 hence y intercept = \(\large{0.50(1.06)^{0}}\) may i know that what r u getting ?

mathslover (mathslover):

remember \(\large{a^0=1}\)

OpenStudy (anonymous):

So would the y intercept be .5?

mathslover (mathslover):

great you are right. . very good

OpenStudy (anonymous):

Thanks! I have one more question if that's okay

mathslover (mathslover):

Sorry i dont know much about growth or decay but since 0.50 is positive hence i will go with Growth

mathslover (mathslover):

by the way : \[\Huge{\mathbb{We}\textbf{lcome}\mathbb{To}\textbf{Open}\mathbb{Study}}\]